OpenStudy (anonymous):

18. Write an equation for the line that is parallel to the given line and passes through the given pointy = 5x + 10; (2, 14) (1 point

OpenStudy (anonymous):

Well for your new line, your slope will be: m=5 Your point will be (2,14). To write an eqn of a line, you need two things: Your slope and your y-intercept, b. To find the y-intercept, you can use your point and your slope to solve for b. y=mx+b 14 = 5(2) +b b = 4 Now all you have to do is plug m and b into y=mx +b, and you have your line! y = 5x +4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

If you have a problem where they give you a point, and state that the line that passes through it is parallel to (another line equation), then you know that each line will have the same slope. If you have your point, and your slope from the other equation then you can find your new line.

OpenStudy (anonymous):

k